In still water, a boat averages 8 miles per hour. It takes the same amount of time to travel 30 miles downstream, with the current, as 18 miles upstream, against the current. What is the rate of the water's current?
Answer by lwsshak3(11628)

n still water, a boat averages 8 miles per hour. It takes the same amount of time to travel 30 miles downstream, with the current, as 18 miles upstream, against the current. What is the rate of the water's current.
..
Let c=speed of current in mph
c+8=speed of boat downstream in mph
c-8=speed of boat upstream in mph
Travel time= distance/speed
30/8+c=18/8-c
240-30c=144+18c
48c=96
c=2 mph
ans
rate of water's current = 2 mph